OpenStudy (anonymous):

when you add: 2x+8x would the answer be 10x^2? or 10x?

OpenStudy (anonymous):

10x

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

10x because when adding terms with variable with same power we just add the coefficients

OpenStudy (anonymous):

10x b'coz x+x=2x but x*x=x^2

OpenStudy (anonymous):

so when you multiply 2x and 8x would it equal to 10x^2?

OpenStudy (anonymous):

ok, i understand, thanks pple

OpenStudy (anonymous):

10x

OpenStudy (anonymous):

it will be 16x^2

myininaya (myininaya):

2 smiley faces+2 smiley face =4 smiley faces not smiley faces squared

OpenStudy (anonymous):

You are not multiplying you are adding, you add the 2 and the 8 to get 10 x

OpenStudy (anonymous):

ok, i got it

OpenStudy (anonymous):

on multiplying u multiply the numbers and add up the powers of the variable so 8x X 2x = 16x^2

OpenStudy (anonymous):

vote if u understand

OpenStudy (anonymous):

Yes pls click on good answer if ur question has been satisfactorily answered

OpenStudy (anonymous):

sorr, im new to this.

OpenStudy (anonymous):

sorry*

OpenStudy (anonymous):

It's like adding two apples and 8 apples, it gives 10 apples, not 10 apples^2 :D

myininaya (myininaya):

anwar i used smiley faces lol

myininaya (myininaya):

and hey

OpenStudy (anonymous):

Haha didn't see it.. You talk a lot about smiley faces lately :D

OpenStudy (anonymous):

and hello!

myininaya (myininaya):

lol should i talk about sad faces?

OpenStudy (anonymous):

I can't say what you should talk about. It's nothing more than a notice :)

myininaya (myininaya):

lol see ya :(

OpenStudy (anonymous):

lol bye *smiley face* :P